Q3. Oil of unknown properties is heated in a shell-and-tube heat exchanger with one shell pass and 20 tube passes. The oil flcan you solve only last 4 questions
Answer #1

a Answer Thi/360k 012 Thi-Teo water Mzoo2 2 kg/s (360-310) = so To=3 lok Tho -3ook Ozc (Tho-Teo 022 300-280 Ti = 280k = 20 OuMaximum rate of heat transfer C Qmax = (min [ Tmax - Tmin ((mm) [Thi-Te 836 [ 360 – 280] Qmax 66880 W 66.88 KW Ane To determi-NTUCIRC) Ez 1-e -NTUCI-3) -NTU (1-0,5) rce 6675 = l-é -NTU (1-05) 1-6.5xe NTO -1.83 Am g over all heat transfer coefficient10 کا The average convection heat transfor Coefficient of the tube outer surface 1 I + 1 hi Ai ho to UA Tvalue of thie Ai innA thumbs up, pls !!